OpenStudy (anonymous):

f(x)=-10x-3 find the slope and y intercept

OpenStudy (anonymous):

point slope form => y=mx+b. This form has 2 parameters. The parameters are (m,b) m is the slope, and b is the y intercept.

OpenStudy (anonymous):

so what's the answer?

jimthompson5910 (jim_thompson5910):

f(x) = -10x-3 is in the form f(x) = mx+b which is basically the same as y = mx+b The general equation y = mx+b has a slope of m and a y intercept of (0, b) In the case of f(x) = -10x-3, we can see that m = -10 and b = -3 So the slope is -10 and the y intercept is (0, -3)

OpenStudy (anonymous):

what is m? what is b? you tell me

OpenStudy (anonymous):

m is -10 of course and b is 3?